OpenStudy (anonymous):

Carbon-14 has a half-life of 5,730 years. How long will it take for 112.5 g of a 120.0-g sample to decay radioactively? 5,730 years 11,460 years 17,190 years 22,920 years

OpenStudy (aaronq):

\(\Large \sf [A]_t=[A]_0*e^{-\frac{ln(2)t}{t_{1/2}}}\) \(\Large \sf (120.0g -112.5 g)=120.0g *e^{-\frac{ln(2)t}{5,730 ~years}}\) \(\sf \Large t=-\dfrac{5,730 ~years*{ln( \frac{(120.0g -112.5 g)}{120.0g})}}{ln(2)}=22386.48 ~years\)